Probability 0.08 0.12 0.30 0.20 0.25 0.05 Telephone Lines in Use 4

Refer to the probability distribution in the table above. Determine the usual minimum and the usual maximum values using the range rule of thumb.
Answer #1
x P(x) xP(x) x2P(x)
0 0.08 0.000 0.000
1 0.12 0.120 0.120
2 0.3 0.600 1.200
3 0.2 0.600 1.800
4 0.25 1.000 4.000
5 0.05 0.250 1.250
total 2.570 8.370
E(x) =μ= ΣxP(x) = 2.5700
E(x2) = Σx2P(x) = 8.3700
Var(x)=σ2 = E(x2)-(E(x))2= 1.765
std deviation=         σ= √σ2 = 1.33

hence

minimum usual value =μ-2*σ = -0.09
maximum usual value =μ+2*σ = 5.23